# Improving LLM Outputs Against Jailbreak Attacks with Expert Model Integration

### v1

@@ III-A Archias - the expert
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=1]fig2.png Performance of LLMs with and without expert model’s output ingested.
The primary dataset for the expert model was derived from three distinct sources: publicly available, synthetic and masked data from Impel, an automotive AI company. By synthetic generation, we refer to the process where we initially crafted prompts and example templates using a team of three people. These templates were then fed into our internally hosted model to generate a synthetic dataset. This dataset is categorized into five categories (see Table I and training samples from the expert model datasets are shown in Appendix, Fig.A):
@@ III-B Benchmark dataset
We manually crafted a total of 150 examples, each designed to challenge the models on various aspects of the five topics in Table  I. The crafting process involved generating inquiries and corresponding choices, and conducting meticulous reviews and revisions to ensure clarity, relevance, and the potential to differentiate between more and less capable models in handling complex linguistic tasks. The challenge was to ensure that the correct answer was not very obvious and required reasoning. With assistance from our colleagues, we achieved this goal. Following the experiment, we invited team members and friends to answer the benchmark questions to gauge human performance. The average score was 88% of the 150 questions answered correctly. To determine human performance, we subsequently invited 50 individuals (30 team members and 20 friends) to answer the same benchmark questions. They were presented with the questions and asked to choose the most appropriate answer - without any additional context or guidance. Additionally, they were instructed not to use any language models or external tools and to determine the answers independently. Upon analyzing the responses, we noticed that certain questions consistently received the same answers from all participants. This prompted us to revise these questions to ensure a more balanced assessment. However, we also discovered that some consistent answers resulted from participant mistakes, not from problems with the questions themselves. Therefore, we decided not to change those questions.
@@ V Conclusion
We will continue to improve defense techniques against prompt injections and explore these concerns, as each improvement in LLM technology is usually accompanied by new dangerous attacks. The effectiveness of this technology opens up new research directions, indicating its potential in a variety of multi-task circumstances. While our benchmark dataset accurately represents the intended themes for testing our models, we realize the need to increase the number of test cases, as jailbreak attempts become increasingly sophisticated. This makes a new research opportunity for us to generalize the approach. Furthermore, we intend to test our automotive LLM with this technology in a production setting, and we will continue to share our experimental findings and latest research initiatives with the community.

### v2

@@ III-A Archias - the expert
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=1]fig2.png Performance of LLMs with and without expert model’s output ingested. The primary dataset for the expert model was derived from three distinct sources: publicly available, synthetic and masked data from Impel, an automotive AI company. By synthetic generation, we refer to the process where we initially crafted prompts and example templates using a team of three prompt engineers. These templates were then fed into our internally hosted model to generate a synthetic dataset. We created 20–30 canonical prompt templates per category (e.g., common price manipulation tactics, malicious intent queries) grounded in publicly known attack patterns and our internal masked data. These templates were manually reviewed and diversified to ensure broad coverage. All generated variations were subsequently subjected to manual quality checks to maintain data integrity. This dataset is categorized into five categories (see Table I and training samples from the expert model datasets are shown in Appendix, Fig.A):
@@
Furthermore, while our current implementation focuses on automotive industry, the same training methodology can be applied to other domains. To extend Archias to a new field, one would need to gather or generate domain-specific inquiry examples (e.g., healthcare, finance), categorize them into relevant classes (e.g., in-domain, out-of-domain), and fine-tune the expert model using the same supervised learning setup described above.
@@ III-B Benchmark dataset
We manually crafted a total of 150 examples, each designed to challenge the models on various aspects of the five topics in Table  I. The crafting process involved generating inquiries and corresponding choices, and conducting meticulous reviews and revisions to ensure clarity, relevance, and the potential to differentiate between more and less capable models in handling complex linguistic tasks. The challenge was to ensure that the correct answer was not very obvious and required reasoning. With assistance from our colleagues (Three of our team’s prompt engineers), we achieved this goal. Following the experiment, we invited team members and friends to answer the benchmark questions to gauge human performance. The average score was 88% of the 150 questions answered correctly. To determine human performance, we subsequently invited 50 individuals: 30 team members (domain knowledge experts) and 20 friends (non-expert participants) to answer the same benchmark questions. They were presented with the questions and asked to choose the most appropriate answer - without any additional context or guidance. Additionally, they were instructed not to use any language models or external tools and to determine the answers independently. Upon analyzing the responses, we noticed that certain questions consistently received the same answers from all participants. This prompted us to revise these questions to ensure a more balanced assessment. However, we also discovered that some consistent answers resulted from participant mistakes, not from problems with the questions themselves. Therefore, we decided not to change those questions. This mixed group reflects both informed and average users, helping us benchmark model performance in realistic deployment settings. We emphasize that the 88% accuracy reflects aggregated human performance and not an expert-only upper bound, providing a practical frame of reference for model comparison.
@@
The complete benchmark dataset, including all examples and corresponding expert-labeled variants, is publicly available on Hugging Face [51]. This allows researchers to reproduce our results and extend evaluations across other models and domains.
@@ V Conclusion
We will continue to improve defense techniques against prompt injections and explore these concerns, as each improvement in LLM technology is usually accompanied by new dangerous attacks. The effectiveness of this technology opens up new research directions, indicating its potential in a variety of multi-task circumstances. While our benchmark dataset accurately represents the intended themes for testing our models, we realize the need to increase the number of test cases, as jailbreak attempts become increasingly sophisticated. Current limitations also include edge-case misclassifications and the need for broader testing under real-world constraints. This makes a new research opportunity for us to generalize the approach. Furthermore, we intend to test our automotive LLM with this technology in a production setting, and we will continue to share our experimental findings and latest research initiatives with the community.